40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1361     Entry time: Thu Mar 5 05:07:09 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Wednesday night locking 
Tonight, I was having a problem with the PO_DC hand-off.
It fails most of the time.
I increased the averaging time for the PD1_DC offset measurement.
I also wrote a script to match the gain of the transmission DC and the PO_DC signals.
This script (/cvs/cds/caltech/scripts/CM/matchPODCGain ) measures the gains of the old (TRX+TRY) and new (PO_DC) signals at 150Hz and returns the optimal value to be put into the input matrix.
cm_step script calls matchPODCGain to determine the matrix element value for the PO_DC signal.

Even with this script, the hand-off was still unreliable.
I checked the AO path loop gain just before the hand off. It looked normal.
Then I realized that the oscilloscope I hooked up to the PO_DC signal using a T-BNC may be introducing some noise into the channel.
So I removed it. Then the PO_DC hand off went well at least once.
The IFO still loses lock at around arm power 10.

I attached time series of the latest lock loss. The second attachment is a zoom of the first one.
This time, there is a glitch in the ETM feedback signals, which is also present in the DARM and CARM and error signals.
I saw this kind of glitches several times today.
Attachment 1: lockLoss5.png  21 kB  | Hide | Hide all
lockLoss5.png
Attachment 2: lockLoss5-zoom.png  23 kB  | Hide | Hide all
lockLoss5-zoom.png
ELOG V3.1.3-